# Exoplanet Detection : A Detailed Analysis

### v1

@@ 3.4 Gravitational Microlensing
The general relativity-based observation method known as gravitational microlensing Perryman (2018). Using gravitational microlensing, scientists found 210 exoplanets by the end of 2023. This technique uses the gravitational potential of a foreground object (the lens) to bend light beams from a distant background object (the source) Bozza et al. (2016) Perryman (2018). Depending on how the source, lens, and observer are aligned, the outcome is a distorted image of the source, sometimes in many images Chen (2023) Beaulieu (2023). The effects of gravitational lensing can be observed, leading to the classification into various domains Jones (2008). Robust lensing, observable at the level of individual objects, is further classified into macro- and microlensing. Whereas microlensing deals with discrete, unresolved multiple visuals, macrolensing generates multiple resolved images, or "arcs." Weak lensing is a different regime that is statistically noticed without a clear resolution of individual objects. Specific circumstances within gravitational lensing are described by the introduction of specialized words. Mesolensing is the term used to characterize the astrometric displacement that accompanied a high-proper motion foreground star, whereas nanolensing is the lensing caused by planetary-mass objects Gaudi (2012). Relative motion between the source, lens, and observer causes time-varying amplification of pictures .
Depending on the characteristics of the radiation source and the lens, this amplification takes place across an assortment of intervals, from hours to years. A more complex light curve may be seen from the background source if the foreground lensing object is gravitationally complex, such a star with planets or a cluster of galaxies Wright & Gaudi (2012). This complexity results from the changing alignment geometry’s time-varying magnification. Significant discoveries have been made using gravitational microlensing Tamanini & Danielski (2019) in exoplanet exploration. In 2004, a 4MJ planet was first clearly detected at approximately 4 astronomical units (au) Perryman (2018). A 5M $\oplus$ planet was then found in 2006. An observation of a two-planet system in 2008 made it possible to quantify orbital motion during the lensing event. In 2015, the lens mass and microlens parallax were measured. In early 2018, evidence for planets flying freely was proposed. These results validate gravitational microlensing as an independent and potent probe of exoplanets, especially over an important mass and orbital radius region Bozza et al. (2016).
@@ 3.5 Astrometry
Detection method Principle Measurements Advantages speciality Drawbacks Equipement
Radial velocity Gravitational Wobble $m_{p},T,a$ large amount of planets can be observe at once, precision large planets, small cool stars its for only planets closer to observer Spectrometer
@@
The exact measurement of the motions and positions of different celestial objects, from bodies of the solar system and stars within our galaxy to, theoretically, entire galaxies and clusters, is known as astrophysicist. Determining stellar parallaxes and proper movements has historically been a major goal of astrometry, as these are essential for comprehending the characteristics of host stars Perryman (2018). In the present development, astrometry plays a new role: it seeks to identify the crosswise element of the displacement of a host star due to an orbiting planet’s gravity. Measures of radial velocity, which are sensitive to the concomitant movement of the photocenter along the line-of-sight, are closely associated with this dynamic manifestation in the sky Saffe et al. (2005).
It has been difficult to achieve high-accuracy astrometry; hipparcos and the fine guidance sensors on the hubble space telescope, for example, achieved accuracies of about 1 milliarcsecond (mas). Still, this degree of accuracy has only just touched the tip of the iceberg in terms of how much star position displacements caused by planets in orbit can be identified. By the end of 2023, astrometry had only found three exoplanets, two of which were 5MJ and 2.26MJ in mass, and one of which was a substellar partner with a mass 28 times that of jupiter (28MJ). These discoveries were recorded in the NASA exoplanet archive. Since the gaia Ranalli et al. (2018) mission was launched in 2013, the field of astrophysicists has seen a dramatic change Panahi et al. (2022). Over a billion stars with magnitudes up to V = 20–21 are being actively measured by gaia, with accuracies of about 20–25 microarcseconds (µas) for stars with a magnitude of 15 Perryman (2018). Our knowledge of planetary systems is predicted to be completely transformed by this astounding increase in precision Hodgkin et al. (2021) Panahi et al. (2022). Several thousand planetary systems with correct planet masses and absolute orbits free of the sin i ambiguity (the uncertainty surrounding the inclination angle of the planetary orbit) should be found thanks to gaia’s large dataset Panahi et al. (2022) Perryman (2018). Moreover, gaia will provide insightful data on planetary systems’ co-planarity, illuminating the configuration and behaviour of several planets within a system wikipedia (2022).
@@ 4 Comparative analysis of detection methods
Count of detected exoplanets by specific method is the proof of its importance. Every method of extrasolar planet detection works on different principles and scientific concepts Perryman (2018) Chen (2023). While transit photometry and radial velocity methods are playing main role for the majority detection. after discovery there are various parameters and properties to know for detecting habitable planet, which is possible through other methods Meunier et al. (2012). hence every detection method have its own advantages and disadvantages Fischer et al. (2015). The transit photometry method coveres wide range of stars but its false positive rate is high Tembhare et al. (2023). transit method works for star-planet systems, which have orbits aligned in such a way that, they can be visible from earth. those exoplanets should move in between host star and earth. Radial velocity method is highly prolific and works for many planets but they must be close to the spectator McArthur et al. (2004). Timing method works for limited materiality, but provide accurate data of exoplanets residing far from earth. Direct imaging also works for few stars but straight forward detection of exoplanets. Large orbital radii planets can be detected by gravitational microlensing, but here measurements alignments can’t be repeated and questionable Hebrard et al. (2018). The majority of precise searches employ multiple methodologies to collect multifaceted information on a certain planet. In particular exoplanet HD 86266 c alternative name TOI-652 c detected by primary transit method in 2020 and gave radius parameter details. Table 3 shows the comparative analysis of detection methods with important parameters.
@@ 5 Space missions for exoplanets
There are several missions are working towards exoplanets and there discoveries. these missions are running by various space agencies like NASA, ESA, JAXA, SETI etc. some missions ended up working and some are currently doing good work. here is the details of some space missions for exoplanets. The landscape of exoplanet exploration has been significantly shaped by a series of pioneering space missions, each contributing unique insights into the vast diversity of planetary systems beyond our solar neighborhood Galicher et al. (2014). Among these missions, TESS (Transiting Exoplanet Survey Satellite) Kane et al. (2020) and GAIA Ranalli et al. (2018) stand as ongoing endeavors,
@@
Space missions for exoplanets
Mission Status
TESS Ongoing
@@
tirelessly scanning the celestial expanse for transiting exoplanets and providing precise astrometric data for over a billion stars, respectively.There are various machine learning techniques Barboza et al. (2020) are used to study and detect exoplanets Jin et al. (2022). Table 4 shows the list of missions working on exoplanet wikipedia (2022) .
Kepler, although completed in 2013, left an enduring legacy by unveiling thousands of exoplanets through the meticulous scrutiny of stellar light curves. Its successor, K2, continued this mission, expanding our understanding of exoplanets until its completion in 2018. Ongoing missions such as CHEOPS (Characterizing Exoplanet Satellite) and JWST (James Webb Space Telescope) Alderson et al. (2023) hold promise for further revelations Agrawal et al. (2015). CHEOPS focuses on characterizing known exoplanets, scrutinizing their atmospheres and compositions, while JWST, a joint venture by NASA, ESA, and CSA, is poised to revolutionize infrared astronomy and deepen our understanding of exoplanet atmospheres Hebrard et al. (2018).

### v2

@@ 3.4 Gravitational Microlensing
The general relativity-based observation method known as gravitational microlensing Perryman (2018). Using gravitational microlensing, scientists found 210 exoplanets by the end of 2023. This technique uses the gravitational potential of a foreground object (the lens) to bend light beams from a distant background object (the source) Bozza $et~{}al$. (2016) Perryman (2018). Depending on how the source, lens, and observer are aligned, the outcome is a distorted image of the source, sometimes in many images Chen (2023) Beaulieu (2023). The effects of gravitational lensing can be observed, leading to the classification into various domains Jones (2008). Robust lensing, observable at the level of individual objects, is further classified into macro- and microlensing. Whereas microlensing deals with discrete, unresolved multiple visuals, macrolensing generates multiple resolved images, or ”arcs.” Weak lensing is a different regime that is statistically noticed without a clear resolution of individual objects. Specific circumstances within gravitational lensing are described by the introduction of specialized words. Mesolensing is the term used to characterize the astrometric displacement that accompanied a high-proper motion foreground star, whereas nanolensing is the lensing caused by planetary-mass objects Gaudi (2012). Relative motion between the source, lens, and observer causes time-varying amplification of pictures .
Depending on the characteristics of the radiation source and the lens, this amplification takes place across an assortment of intervals, from hours to years. A more complex light curve may be seen from the background source if the foreground lensing object is gravitationally complex, such a star with planets or a cluster of galaxies Wright & Gaudi (2012). This complexity results from the changing alignment geometry’s time-varying magnification. Significant discoveries have been made using gravitational microlensing Tamanini & Danielski (2019) in exoplanet exploration. In 2004, a 4MJ planet was first clearly detected at approximately 4 astronomical units (au) Perryman (2018). A 5M $\oplus$ planet was then found in 2006. An observation of a two-planet system in 2008 made it possible to quantify orbital motion during the lensing event. In 2015, the lens mass and microlens parallax were measured. In early 2018, evidence for planets flying freely was proposed. These results validate gravitational microlensing as an independent and potent probe of exoplanets, especially over an important mass and orbital radius region Bozza $et~{}al$. (2016).
@@ 3.5 Astrometry
[…]
@@
The exact measurement of the motions and positions of different celestial objects, from bodies of the solar system and stars within our galaxy to, theoretically, entire galaxies and clusters, is known as astrophysicist. Determining stellar parallaxes and proper movements has historically been a major goal of astrometry, as these are essential for comprehending the characteristics of host stars Perryman (2018). In the present development, astrometry plays a new role: it seeks to identify the crosswise element of the displacement of a host star due to an orbiting planet’s gravity. Measures of radial velocity, which are sensitive to the concomitant movement of the photocenter along the line-of-sight, are closely associated with this dynamic manifestation in the sky Saffe $et~{}al$. (2005).
It has been difficult to achieve high-accuracy astrometry; hipparcos and the fine guidance sensors on the hubble space telescope, for example, achieved accuracies of about 1 milliarcsecond (mas). Still, this degree of accuracy has only just touched the tip of the iceberg in terms of how much star position displacements caused by planets in orbit can be identified. By the end of 2023, astrometry had only found three exoplanets, two of which were 5MJ and 2.26MJ in mass, and one of which was a substellar partner with a mass 28 times that of jupiter (28MJ). These discoveries were recorded in the NASA exoplanet archive. Since the gaia Ranalli $et~{}al$. (2018) mission was launched in 2013, the field of astrophysicists has seen a dramatic change Panahi $et~{}al$. (2022). Over a billion stars with magnitudes up to V = 20–21 are being actively measured by gaia, with accuracies of about 20–25 microarcseconds (µas) for stars with a magnitude of 15 Perryman (2018). Our knowledge of planetary systems is predicted to be completely transformed by this astounding increase in precision Hodgkin $et~{}al$. (2021) Panahi $et~{}al$. (2022). Several thousand planetary systems with correct planet masses and absolute orbits free of the sin i ambiguity (the uncertainty surrounding the inclination angle of the planetary orbit) should be found thanks to gaia’s large dataset Panahi $et~{}al$. (2022) Perryman (2018). Moreover, gaia will provide insightful data on planetary systems’ co-planarity, illuminating the configuration and behaviour of several planets within a system wikipedia (2022).
@@ 4 Comparative analysis of detection methods
Count of detected exoplanets by specific method is the proof of its importance. Every method of extrasolar planet detection works on different principles and scientific concepts Perryman (2018) Chen (2023). While transit photometry and radial velocity methods are playing main role for the majority detection. after discovery there are various parameters and properties to know for detecting habitable planet, which is possible through other methods Meunier $et~{}al$. (2012). hence every detection method have its own advantages and disadvantages Fischer $et~{}al$. (2015). The transit photometry method coveres wide range of stars but its false positive rate is high Tembhare $et~{}al$. (2023). transit method works for star-planet systems, which have orbits aligned in such a way that, they can be visible from earth. those exoplanets should move in between host star and earth. Radial velocity method is highly prolific and works for many planets but they must be close to the spectator McArthur $et~{}al$. (2004). Timing method works for limited materiality, but provide accurate data of exoplanets residing far from earth. Direct imaging also works for few stars but straight forward detection of exoplanets. Large orbital radii planets can be detected by gravitational microlensing, but here measurements alignments can’t be repeated and questionable Hebrard $et~{}al$. (2018). The majority of precise searches employ multiple methodologies to collect multifaceted information on a certain planet. In particular exoplanet HD 86266 c alternative name TOI-652 c detected by primary transit method in 2020 and gave radius parameter details. Table 3 shows the comparative analysis of detection methods with important parameters.
@@ 5 Space missions for exoplanets
There are several missions are working towards exoplanets and there discoveries. these missions are running by various space agencies like NASA, ESA, JAXA, SETI etc. some missions ended up working and some are currently doing good work. here is the details of some space missions for exoplanets. The landscape of exoplanet exploration has been significantly shaped by a series of pioneering space missions, each contributing unique insights into the vast diversity of planetary systems beyond our solar neighborhood Galicher $et~{}al$. (2014). Among these missions, TESS (Transiting Exoplanet Survey Satellite) Kane $et~{}al$. (2020) and GAIA Ranalli $et~{}al$. (2018) stand as ongoing endeavors,
[…]
@@
[…]
@@
tirelessly scanning the celestial expanse for transiting exoplanets and providing precise astrometric data for over a billion stars, respectively.There are various machine learning techniques Barboza $et~{}al$. (2020) are used to study and detect exoplanets Jin $et~{}al$. (2022). Table 4 shows the list of missions working on exoplanet wikipedia (2022) .
Kepler, although completed in 2013, left an enduring legacy by unveiling thousands of exoplanets through the meticulous scrutiny of stellar light curves. Its successor, K2, continued this mission, expanding our understanding of exoplanets until its completion in 2018. Ongoing missions such as CHEOPS (Characterizing Exoplanet Satellite) and JWST (James Webb Space Telescope) Alderson $et~{}al$. (2023) hold promise for further revelations Agrawal $et~{}al$. (2015). CHEOPS focuses on characterizing known exoplanets, scrutinizing their atmospheres and compositions, while JWST, a joint venture by NASA, ESA, and CSA, is poised to revolutionize infrared astronomy and deepen our understanding of exoplanet atmospheres Hebrard $et~{}al$. (2018).